Shift report 20000722 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Jul 22 2000 - 16:15:09 EDT

  • Next message: http account: "Shift report 20000723 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Jul 22 16:15:08 2000, Shift: 08:00-16:00
    Supervisor: PC, with: DO
    Log book#3 pages 44 - 46
    
    Shift Summary:
    ------------------------------------------------------------
    Tried to ramp beam a few times.
    Beam intensities at injection energies are ~ 2*10^9 for both
    beams.
    Blue beam seems stable at injection energy. Loose half 
    when ramping and seems stable afterwards.
    Yellow beam seems much less stable at injection energies
    and is usually lost when ramping ( It survived once
    at 44*10^6).
    
    When doing shift check we noticed that connection to
    tsir was refused ( Means no talk to H1,H2, C1, and TOFW).
    TPCs are still reachable although they are on tsir.
    KO and FV were informed. KO talked to MCR and was granted an
    access if needed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Jul 22 2000 - 16:18:22 EDT